The density of chloroform, a widely used organic solvent, is 1.4832 g/mL. How many milliliters would you use if you wanted 112.5 g of chloroform?
1
Expert's answer
2021-09-28T01:03:53-0400

 1.4832/1=112.5/v

where v is the volume.

Cross-multiply:

1.4832v= 112.5

so v= 112.5/1.4832

=75.85ml